How much alcohol is required to preserve 2 L of a syrup containing 1000 g sucrose? No API is included in the syrup.  
Answer #1

Answer: Syrup USP, having a specific gravity of 1.313 and a concentration of 85% w/v is a 65% w/w solution. This 65% by weight is the minimum amount of sucrose which will preserve neutral syrup. To calculate the preservatives one may assume that free water is preserved by 18% alcohol. But here the syrup is 2 L while it conatines 1000g of sucrose. So, at this concentration alcohol concentration will be 156 ml.

calculations:

In Syrup, USP 850 g sucrose occupies an apparent volume of 550 ml; so each gram of sucrose will occupy 550/850 or 0.647 ml. If the 850 g sucrose preserves 450 ml of water, then each gram of sucrose will preserve 450/850 = 0.53 ml of water.

Volume preserved by sucrose= 1000gm * 0.53 ml= 530ml

Volume occupies by sucrose= 1000gm * 0.647ml= 647 ml

Free water equivalent = 2000-647- 530 = 823 ml

Volume of alcohol required to preserve the syrup = 823 * 18% = 148.14 ml

so, 148.14 ml alcohol divided by 95% = 155.93 ml ( 156 ml)